Archie Co. purchased a framing machine for $51,000 on January 1, 2021. The machine is expected to have a four-year life, with a residual value of $7,000 at the end of four years.

Using the straight-line method, depreciation for 2022 and book value at December 31, 2022, would be:

Multiple Choice

  • $11,000 and $22,000 respectively.

  • $11,000 and $29,000 respectively.

  • $12,750 and $18,500 respectively.

  • $12,750 and $25,500 respectively.
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(51,000-7000)/4=$11000

Hence depreciation for 2022=11000

Book value at December 31, 2022=Cost-Accumulated Depreciation

=51,000-(11000*2)=$29,000

Hence the correct option is :

$11,000 and $29,000 respectively.